difference quotient f(x)= -2x^2
Answer by stanbon(75887)

difference quotient
f(x)= -2x^2
-----
f(x+h) = -2(x+h)^2 = -2x^2 -2(2hx) -2(h^2) = -2x^2 - 4hx -2h^2
----------------
Quotient::
[f(x+h)-f(x)]/h = [-4hx-2h^2]/h = -4x - 2h
-----------------
Cheers,
Stan H.
----------